P2 Establish local collection centers throughout the region to decrease the time it takes to convert...

P2
Establish local collection centers throughout the region to decrease the time it takes to convert credit payments that are mailed in by check to cash. It is estimated that establishing these collection centers would reduce the average collection time by 2 days (from 5 days to 3 days).

  1. If the company currently averages $20,000 in collections per day, how many dollars will this suggested cash management system frees up?
  2. If all freed up dollars would be used to pay down debt that has an interest rate of 8%, how much money could be saved each year in interest expense?

Do the numbers suggest that this new system should be implemented if its total annual cost is $5200? Explain.
Answer #1

1: Amount freed up = No of days * amount per day

=2*20000 = $40,000

2: Amount saved in interest = 40000*8% = 3200

3: No. If the total annual cost is $5200, the cost will be more than the savings. Hence the system should not be implemented.